# Pelican-VL 1.0: A Foundation Brain Model for Embodied Intelligence

### v1

@@ Abstract
Figure 1: Performance comparison of Pelican-VL1.0. (Left) Comparison against models with $\leq$100B parameters. The shaded (pink) region highlights the performance gain over our baseline. (Right) Comparison against models with ¿100B parameters, including leading open-source and proprietary models, where our model also demonstrates SOTA performance.
@@ 3.1 Data for Training
To advance the capability of Pelican-VL 1.0 in perception, reasoning, and planning in real-world scenarios grounded in physical knowledge, we built a diverse data pool totaling 4B tokens. This pool contains 231M images and 29k hours of video, encompassing 231M open-ended question-answer (QA) pairs, 9M grounding annotations (bounding boxes or keypoints), and 2M multiple-choice questions (MCQs). From this curated data, we sampled 1.3M instances for SFT stages and 0.5M for RL, and categorizing the data into four fundamental capability areas to specifically address challenges in embodied AI:
@@
  • Decision Making and Task Planning. This category targets high-level cognitive abilities for robotics. The data is designed to support decision-making in dynamic contexts (e.g., COSMOS) and to translate language instructions into concrete actions (e.g., Where2Place). This includes specific downstream tasks such as determining vacant spaces for placement (Vacant QA) and executing object placement plans (e.g., RefSpatial).
@@ 3.2 Metaloop Data Selection
To address data scarcity in embodied AI, we explore leveraging natural world data (e.g., YouTube videos) for model improvement. We select the SpatialVID wang2025spatialvid dataset but discard its existing expert-derived annotations (e.g., camera pose, spatial captions) to focus on self-evolution from raw video. We use a general-purpose VLM, Qwen3VL-Plus qwen3-vl , to generate 24 spatial QA pairs for each of 75k videos (e.g., ¡How many green cars in this video?¿, ¡The man with the blue jacket and red shirt, who is closer to the camera?¿). We then filter this data using InternVL3.5-38B internvl3.5 , which infers answers for each question twice. A QA pair is kept if InternVL3.5’s answer matches the Qwen3VL-Plus answer. If InternVL3.5’s two inferences are identical to each other but differ from the original, we replace the original answer with the new consensus answer. This refinement process yields 14k QAs.
@@ 4.3.1 Benchmark Performance
Table 2: Overall performance comparison on benchmarks. Bold and underlined numbers indicate the best and second-best results, respectively. An asterisk (*) denotes results reported from official sources. Yellow cells mark our proposed Pelican-VL 1.0 models.
@@
Spatial-Physical EgoSchema 57.5 61.2 66.5 71.4 68.2 70.5 69.0 70.9 73.3 79.3
@@
(b) Performance of Models with ¿100B Parameters.
Category Benchmark Qwen3-VL Qwen3-VL InternVL3.5 Qwen3-VL-Plus GPT-5 Gemini2.5-Flash GPT-4o Pelican-VL
235B-A22B-Thinking 235B-A22B-Instruct 241B-A28B proprietary proprietary proprietary proprietary 72B
Common MVBench 74.5 76.1 78.4 76.4 66.9 65.2 64.4 69.7
Spatial-Physical EgoSchema 72.1 77.7 72.9 78.0 68.8 61.8 69.2 79.3
BLINK 67.4 68.6 61.4* 69.3 62.8* 62.6 64.2 60.3
PhyX 85.3 63.2 62.6 63.9 82.7 77.7 41.2 86.4
OmniSpatial 53.8 52.1 53.2 52.9 56.2 43.6 41.7 49.6
Where2Place 52.2 40.0 35.1 46.3 29.6 35.1 20.3 64.0
RoboSpatial 62.0 58.0 51.1 59.1 43.6* 59.9 46.9 61.1
EmbSpatiaBench 83.4 82.4 79.1 82.6 75.1* 74.6 71.9 76.6
RefSpatialBench 39.4 20.6 23.6 19.9 23.1 35.4 9.9 49.5
CoT-Reasoning ERQA 48.0 47.5 47.3 48.0 42.0* 49.8 37.8 43.0
COSMOS 62.3 60.4 56.4 53.0 61.8 30.3 52.4 68.5
VSI-Bench 60.9 62.6* 67.0 61.2 48.2 46.4 34.0 57.3
Average 63.4 59.1 57.4 59.2 55.1 53.5 46.2 63.8
@@ 6.1 Embodied Foundation Model
The central challenge in embodied AI is the creation of a generalist Embodied Foundation Model (EFM)—a scalable robotic brain that unifies perception, reasoning, and action. As noted in our introduction, the field has largely bifurcated into two primary strategies to achieve this. The first strategy is heterogeneous data scaling, where major labs demonstrate high performance by adapting massive, pre-trained VLMs through data pyramid scaling. This data-scaling paradigm is exemplified by flagship efforts such as Google’s Gemini Robotics abdolmaleki2025gemini , which leverages its powerful VLM for complex reasoning in physical tasks; NVIDIA’s GR00T N1 bjorck2025gr00t & Cosmos-Reason1 azzolini2025cosmos , open models explicitly designed for generalist humanoid control; ByteDance’s GR-3 cheang2025gr , which integrates web-scale knowledge; Physical Intelligence’s $\pi_{0.5}$ intelligence2504pi0 , which uses novel flow-based architectures for open-world generalization; RoboBrain Ji2025RoboBrain and others figure2025helix . While these approaches have pushed the boundaries of performance, they often rely on the sheer scale of the data, lacking a sophisticated algorithmic framework to manage, refine, and consolidate learning from such vast, heterogeneous, and often noisy sources. The second strategy is architectural refinement, which proposes more decoupled, layered systems. This approach, seen in models like Figure’s Helix and Wall-OSS zhai2025igniting , typically uses a large VLM as a high-level “brain” for reasoning and task planning, which in turn commands a separate, smaller, reactive policy for low-level motor control. While architecturally elegant, this paradigm often operates on smaller, more specialized datasets, thus lacking the sheer scale and data diversity required for true, open-world generalization.

### v2

@@ Abstract
Figure 1: Performance comparison of Pelican-VL1.0. (Left) Comparison against models with $\leq$100B parameters. The shaded (pink) region highlights the performance gain over our baseline. (Right) Comparison against models with >>100B parameters, including leading open-source and proprietary models, where our model also demonstrates SOTA performance.
@@ 3.1 Data for Training
To advance the capability of Pelican-VL 1.0 in perception, reasoning, and planning in real-world scenarios grounded in physical knowledge, we built a diverse data pool. This pool contains 231M images and 29k hours of video, encompassing 231M open-ended question-answer (QA) pairs, 9M grounding annotations (bounding boxes or keypoints), and 2M multiple-choice questions (MCQs) From this curated data, we sampled 1.3M instances for SFT stages and 0.5M for RL, amounting to 4B training tokens, and categorizing the data into four fundamental capability areas to specifically address challenges in embodied AI:
@@
  • Decision Making and Task Planning. This category targets high-level cognitive abilities for robotics. The data is designed to support decision-making in dynamic contexts (e.g., COSMOS) and to translate language instructions into concrete actions (e.g., Where2Place). This includes specific downstream tasks such as determining vacant spaces for placement (e.g., Vacant QA) and executing object placement plans (e.g., RefSpatial).
@@ 3.2 Metaloop Data Selection
To address data scarcity in embodied AI, we explore leveraging natural world data (e.g., YouTube videos) for model improvement. We select the SpatialVID wang2025spatialvid dataset but discard its existing expert-derived annotations (e.g., camera pose, spatial captions) to focus on self-evolution from raw video. We use a general-purpose VLM, Qwen3VL-Plus qwen3-vl, to generate 24 spatial QA pairs for each of 75k videos (e.g., <How many green cars in this video? >, <The man with the blue jacket and red shirt, who is closer to the camera?>). We then filter this data using InternVL3.5-38B internvl3.5, which infers answers for each question twice. A QA pair is kept if InternVL3.5’s answer matches the Qwen3VL-Plus answer. If InternVL3.5’s two inferences are identical to each other but differ from the original, we replace the original answer with the new consensus answer. This refinement process yields 14k QAs.
@@ 4.3.1 Benchmark Performance
Table 2: Overall performance comparison on benchmarks. Bold and underlined numbers indicate the best and second-best results, respectively. A dagger ($\dagger$) marks results differing from official reports or unusually low, possibly as official evaluations used model-specific prompts and the models are prompt-sensitive, while our results are obtained under a unified protocol for fair comparison. An asterisk (*) denotes results reported from official sources. Yellow cells mark our proposed Pelican-VL 1.0 models.
@@
Spatial-Physical RoboSpatial 44.3 48.0 48.6 47.4 57.4 44.3 56.3 47.7 57.5 61.1
@@
(b) Performance of Models with >> 100B Parameters.
Category Benchmark Qwen3-VL Qwen3-VL InternVL3.5 Qwen3-VL-Plus GPT-5 GPT-5-Mini Gemini2.5-Flash GPT-4o Pelican-VL
235B-A22B-Thinking 235B-A22B-Instruct 241B-A28B proprietary proprietary proprietary proprietary proprietary 72B
Common MVBench 74.5 76.1 78.4 76.4 73.1 66.9 65.2 64.4 69.7
Spatial-Physical RoboSpatial 62.0 58.0 51.1 59.1 53.4 50.6 59.9 46.9 61.1
BLINK 67.4 68.6 61.4* 69.3 69.9 67.4 62.6 64.2 60.3
PhyX 85.3 63.2 62.6 63.9 83.6 82.5 77.7 41.2 86.4
OmniSpatial 53.8 52.1 53.2 52.9 57.6 56.2 43.6 41.7 49.6
Where2Place 52.2 40.0 35.1 46.3 38.6 31.5 35.1 20.3 64.0
EgoSchema 72.1 77.7 72.9 78.0 73.7 68.8 61.8 69.2 79.3
EmbSpatialBench 83.4 82.4 79.1 82.6 82.3 79.3 74.6 71.9 76.6
RefSpatialBench 39.4 20.6 23.6 19.9 21.6 13.9 35.4 9.9 49.5
CoT-Reasoning ERQA 48.0 47.5 47.3 48.0 60.0 53.5 49.8 37.8 43.0
COSMOS 62.3 60.4 56.4 53.0 64.8 61.8 30.3 52.4 68.5
VSI-Bench 60.9 62.6* 67.0 61.2 56.2 48.2 46.4 34.0 57.3
Average 63.4 59.1 57.4 59.2 61.2 56.7 53.5 46.2 63.8
@@ 6.1 Embodied Foundation Model
The central challenge in embodied AI is the creation of a generalist Embodied Foundation Model—a scalable robotic brain that unifies perception, reasoning, and action. By its very definition, such a brain is innately multimodal. This pursuit has therefore been directly catalyzed by the profound advancements in large VLMs zhang2024vision; liu2025nexus; team2023gemini; wen2025infmasking, which provide the foundational architecture for this synthesis. As noted in our introduction, the field has largely bifurcated into two primary strategies to achieve this. The first strategy is heterogeneous data scaling, where major labs demonstrate high performance by adapting massive, pre-trained VLMs through data pyramid scaling. This data-scaling paradigm is exemplified by flagship efforts such as Google’s Gemini Robotics abdolmaleki2025gemini, which leverages its powerful VLM for complex reasoning in physical tasks; NVIDIA’s GR00T N1 bjorck2025gr00t & Cosmos-Reason1 azzolini2025cosmos, open models explicitly designed for generalist humanoid control; ByteDance’s GR-3 cheang2025gr, which integrates web-scale knowledge; Physical Intelligence’s $\pi_{0.5}$ intelligence2504pi0, which uses novel flow-based architectures for open-world generalization; RoboBrain Ji2025RoboBrain and others figure2025helix; fan2025xr. While these approaches have pushed the boundaries of performance, they often rely on the sheer scale of the data, lacking a sophisticated algorithmic framework to manage, refine, and consolidate learning from such vast, heterogeneous, and often noisy sources. The second strategy is architectural refinement, which proposes more decoupled, layered systems. This approach, seen in models like Figure’s Helix and Wall-OSS zhai2025igniting, typically uses a large VLM as a high-level “brain” for reasoning and task planning, which in turn commands a separate, smaller, reactive policy for low-level motor control. While architecturally elegant, this paradigm often operates on smaller, more specialized datasets, thus lacking the sheer scale and data diversity required for true, open-world generalization.
@@ 7. Discussion and Future Work
8. Contributions
Our contributors are organized based on their roles and magnitude of contribution. The specific roles are as follows:
8.1 Core Contributors
  • Yi Zhang, Che Liu, and Xiancong Ren
8.2 Contributors
  • Hanchu Ni, Shuai Zhang, Zeyuan Ding, Yingji Zhang, Jiayu Hu, Haozhe Shan, Zhenwei Niu, Zhaoyang Liu, Shuang Liu, Yue Zhao, Junbo Qi, Yan Bai, Qinfan Zhang, Dengjie Li, Yidong Wang, Jiachen Luo, Zenglin Xu, Bin Shen, and Qifan Wang
  • Baidu AI Cloud Baige Team (aihc@baidu.com)
8.3 Tech Lead
  • Yong Dai
8.4 Corresponding Authors
  • Jian Tang and Xiaozhu Ju
8.5 Acknowledgement
We would like to express our sincere gratitude to all our colleagues at X-Humanoid for their invaluable support and insightful discussions throughout this ambitious project. We are particularly grateful to Wanxin Tian from the Agent team; his persistent feedback and deep insights into the agent’s iterative development were instrumental in shaping our research. Our heartfelt thanks also go to Fangzhe Yan, Wenwen Han, Zexin Yang, Ouyang Li and the entire data annotation team. Their diligent and meticulous work in processing and labeling our petabyte-scale dataset was foundational to the success of Pelican-VL 1.0.
9. Appendix
9.1 Grasp Contact Model and Adaptive Grasp
Human can sense the friction information on contact surface based on the slip state of the fingertips to control the grasping force [johansson1987signals]. Therefore, slip detection is important for grasp performance. A promising approach for slip detection is to quantify the incipient slip which means to detect the distance from the macro slip (where the contact surfaces of two objects have completely slid with respect to each other) [sui2023novel]. Sui [sui2023novel] and Li [li2025learning] proposed a contact factor based on the vision-based tactile sensor (VBTS) to quantify incipient slip and enable reactive grasp control. In this framework, we proposed a stochastic model based on the contact factor and realized a synchronized online friction estimation and adaptive grasp control.
The contact coefficient is defined as:
$cf=1-\frac{F_{\text{MER},t}}{\mu\cdot F_{\text{MER},n}}$ (15)
where $F_{\text{MER},t}$ and $F_{\text{MER},n}$ represent the integral of tangential and normal force magnitudes over the contact surface, respectively. The contact coefficient, denoted as $cf$, is bounded between 0 and 1. A value of $cf$ close to 0 indicates that the contact state is to macro slip, while values closer to 1 reflect greater contact stability. This metric provides a quantitative measure of grasp security and can be directly incorporated into adaptive grasping control.
In the above contact model, the friction coefficient $\mu$ is essential to be given as a prior. However, this prerequisite is often infeasible to get for arbitrary objects. This limitation has led to solutions misaligned with the goal of human-like adaptability. For instance, Li et al. [li2025learning] employ a two-stage estimation and control process that is prohibitively slow. While the online estimation method in [sui2023novel] is a conceptual advance, its multi-second convergence time remains significantly slower than human performance and is insufficient for real-time applications.
A fundamental limitation underlying above robotic grasping strategies is the decoupling of friction estimation from grasp control. In contrast, human grasping demonstrates a tightly synchronized process where haptic feedback and motor control are seamlessly integrated to achieve reliable manipulation [johansson2009coding]. To endow robots with similar adaptive capabilities, two primary challenges must be addressed: (1) achieving rapid and robust online friction estimation and (2) developing a control framework that seamlessly integrates this estimation in real time. Therefore, we frist proposed a synchronized strategy to achieve the continuous integration of sensing and control that characterizes human sensorimotor coordination.
In contrast to treating friction as a fixed, deterministic parameter, we model it as a stochastic variable. Therefore, we define the system state as $\mu_{t}$. Its state evolution is modeled as a random walk process:
$\mu_{t}$ $=\mu_{t-1}+\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}(0,\sigma_{p}^{2})$ (16)
$\mu_{t}$ $=\text{min}(\text{max}(\mu_{t}\mu_{min}),\mu_{max})$ (17)
where $\sigma_{p}^{2}$ represents process noise variance. Here, $\mu_{min}$ and $\mu_{max}$ are the physical constrains boundary. The state of interest is referred to as the posterior state, $p(\mu_{t}|z_{1:t})$, where $z_{1:t}={z_{1},z_{2},\ldots,z_{t}}$ represents the sequence of all observations from the start of the grasp until time $t$. Then the particle filter is used to estimate the friction coefficient $\mu_{t}$ in the real time.
$p(\mu_{0:t}\mid z_{1:t})\approx\sum_{i=1}^{M}w_{t}^{i}\delta(\mu_{0:t}-\mu_{0:t}^{i})$ (18)
where $\{w_{t}^{i},\mu_{0:t}^{i}\}_{i=1}^{M}$ is the set of $M$ weighted samples. The weight $w_{t}^{i}\in(0,1]$ signifies the relative importance of the sample $\mu_{0:t}^{i}$, with $\sum_{i=1}^{M}w_{t}^{i}=1$. Samples with higher weights correspond to more probable estimates of the true state sequence. Here, $\delta(\cdot)$ denotes the Dirac delta function, which is zero everywhere except at the origin and integrates to one.
The process of grasping, lifting, and replacing an object can be decomposed into seven distinct phases: reach, load, lift, hold, replace, unload, and release [johansson2009coding]. During the reach phase, the objective is to establish contact between the object and the tactile sensor-equipped gripper. The gripper closes gradually until a contact is detected, which is determined by a threshold on the resultant normal force: $F_{n}>F_{n}^{\text{init}}$, where $F_{n}^{\text{init}}$ is the predicted initial force got from Pelican-VL 1.0.
Subsequently, during the load phase, the friction coefficient estimation enables the online update of the contact coefficient $cf_{t}$ via (15), utilizing the measured forces. A proportional control law adjusts the gripper position $x_{g}$ to drive $cf_{t}$ towards a predefined target $cf_{\text{target}}$, which defines the desired safety margin against slip. The gripper command is given as following, and $K_{p}$ is a proportional gain.
$\Delta x_{g}(t)=K_{p}\cdot(cf_{\text{target}}-cf_{t})$ (19)
It is noteworthy that the friction estimation module and the adaptive grasp controller form a closed-loop system: the controller utilizes the current best estimate to adjust the gripper position, while new tactile feedback resulting from this adjustment is continuously incorporated to refine the estimation. This reciprocal interaction establishes a highly responsive and disturbance-resistant sensorimotor cycle.
9.2 Infrastructure as Enabler: Toward a Scalable and Expansive Roadmap
This section details our roadmap for operationalizing our theory, which is built upon two pillars of adjusted applications: a parallelization strategy (Context Parallelism) to scale context length and a custom training framework to conquer the challenges of training a 72B-scale multi-modal model with RL.
Context Parallelism: Scaling Multi-modal Context and Efficiency.
While DAPO optimizes the learning from each data point, the increasing complexity of multi-modal, long-horizon tasks necessitates processing ever-larger contexts, including long dialogue histories, sequential visual observations, and action histories. To overcome the memory bottlenecks of traditional data parallelism, we adopt and optimize Context Parallelism.
Core Mechanism: Context Parallelism addresses this challenge by distributing the input context (tokens, image patches) across multiple GPUs for a single forward pass. Split Context: The input sequence is partitioned into smaller chunks, with each chunk assigned to a different GPU. Parallel Attention: Each GPU computes the attention mechanism for its local chunk while communicating intermediate Key-Value (KV) states with other GPUs to maintain a global context. This allows the model’s context window to scale linearly with the number of available devices, enabling the processing of vast temporal and spatial information crucial for embodied agents.
The Apex Challenge: Enabling RLHF on a 72B Multi-modal Model
. The ultimate ambition of our work is to build a 72-billion parameter multi-modal hybrid model. A significant gap in the current open-source ecosystem is the lack of robust frameworks capable of handling RL from Human Feedback (RLHF) for multi-modal models at such a scale.
Our Solution: To solve this, we adapt and significantly extend the VERL framework. Our primary contribution is the development of a novel training pipeline that enables a 72B-scale model to be trained with our preference learning algorithms on complex, mixed-modality data. Mixed-Batch Multi-modal Training: The cornerstone of our innovation is a sophisticated batching and data loading mechanism that allows a single training batch to concurrently contain text-only, image-text, and video-text data. This heterogeneous batching capability is critical for efficient multi-modal learning. Scaling GRPO to 72B: By leveraging our modified VERL framework and Context Parallelism, we successfully implement the GRPO algorithm at the 72B scale on these mixed-modality batches. This allows the model to learn from rich, ranked preference signals that span across different data types simultaneously, a capability that, to our knowledge, has not been demonstrated in open-source models of this magnitude. Conquering Engineering Hurdles: This achievement required systematically solving a cascade of engineering challenges, including: Extreme Parallelism: Integrating a hybrid of data, pipeline, and tensor parallelism to manage the model’s immense size. Multi-modal Fusion at Scale: Designing robust cross-attention mechanisms that effectively fuse information from disparate modalities without succumbing to training instability. Inference Optimization: Developing quantization and deployment strategies to make such a large model usable in practical applications.
By systematically addressing these infrastructural, algorithmic, and architectural challenges, our roadmap unlocks the full potential of large multi-modal models, providing a concrete and scalable path towards truly intelligent and capable embodied AI. Our framework is built upon the unification of SFT and RL through the lens of preference learning, implemented via a multi-stage and iterative training pipeline.
9.3 Reward Design for Embodied Tasks
To accommodate the diverse modalities and objectives of embodied tasks, we design task-specific reward functions, all unified under a common interface for the RL framework.
(1) Format Reward.
The format reward ensures compliance with the required output pattern, enforcing consistent reasoning syntax without constraining intermediate “thinking” steps:
$r_{\text{format}}=\begin{cases}1,&\text{if the output contains a valid }\\ 0,&\text{otherwise.}\end{cases}$ (20)
This soft structural constraint stabilizes learning and ensures that generated outputs are parsable for downstream evaluation.
(2) Choice Reward.
For standard classification or multiple-choice questions, correctness is directly evaluated as:
$r_{\text{choice}}=\mathbb{1}(\text{pred}=\text{label})$ (21)
which encourages precise logical discrimination and serves as a baseline for assessing embodied question-answering consistency.
(3) Numeric Reward (Counting / Distance).
For tasks yielding quantitative outcomes—such as counting visible objects or estimating spatial distances—we employ a continuous reward reflecting numerical accuracy:
$r_{\text{numeric}}=\exp(-\alpha|\hat{y}-y|)$ (22)
where $\hat{y}$ is the predicted numeric value, $y$ is the ground truth, and $\alpha$ controls sensitivity to error. This design rewards approximate correctness and provides smoother gradients for regression-like embodied reasoning.
(4) Affordance Reward.
For affordance-centric tasks (e.g., determining which object can be grasped, pushed, or opened), we define:
$r_{\text{affordance}}=\text{sim}(\text{pred},\text{ground truth})$ (23)
where $\text{sim}(\cdot)$ can be instantiated as semantic matching or embedding-based alignment between the predicted and target affordance categories. This reward encourages the model to ground perception in actionable affordance semantics.
9.4 Benhmark Analysis
Table 4: Task examples for each dimension in our benchmark.
Primary dimensions Secondary dimensions Example
Physical & Causal Physical feasibility Can the tissue box fit behind the chair? — Yes.
Reasoning Stability and support Will the stacked cups remain stable if the top one is pushed slightly? — No.
Intuitive physics and collisions If the rolling ball hits the block, will it fall over? — Yes.
Perception & Object Click or coordinate prediction Please point out the orange box. — (x, y) = (0.43, 0.72).
Grounding Bounding box detection Identify the bounding box of the cup on the table. — (x1, y1, x2, y2).
Grasp point Where should the robot grasp the handle of the kettle? — At the handle center.
Quantitative & Numerical Counting How many apples are on the table? — Three.
Reasoning Continuous estimation What is the distance between the box and the wall? — About 0.5 meters.
Relative magnitude Which container holds more water? — The right one.
Spatial & Geometric Spatial relations Can the speaker fit left of the cup? — Yes.
Reasoning Egocentric vs allocentric frames From the robot’s view, which object is on its right? — The red box.
Topological relations Which object is inside the basket? — The towel.
Temporal & Sequential Temporal order What happened before the man picked up the ball? — He bent down.
Reasoning State transition and action effects After the cup is tilted, what happens to the water? — It spills.
Motion and sequence reasoning What is the person doing after opening the door? — Walking inside.
Affordance & Function Graspability Which part of the screwdriver should be grasped? — The handle.
Reasoning Openability or pourability Can the bottle be opened by twisting the cap? — Yes.
Tool-use reasoning What tool can be used to cut the rope? — Scissors.
Manipulation feasibility Is it possible to rotate the knob with one hand? — Yes.
Multi-Object & Scene Multi-object relations Which object is partially hidden behind the monitor? — The keyboard.
Consistency Scene-state consistency Is the cup in the same place across both frames? — Yes.
Spatial–language consistency Does the description match the image scene? — No.
Scene & Action Scene types What type of environment is this image captured in? — Kitchen.
Understanding Action recognition What is the person doing in the video? — Pouring water.
State recognition Is the fridge door open or closed? — Open.
Pose, gesture, direction What direction is the person pointing to? — Left.
Decision & Task Next-action prediction What should the robot do after picking up the cup? — Place it on the tray.
Planning Sequential plan List the steps to make a cup of tea. — Boil water → Add tea → Pour → Serve.
Sub-goal decomposition What are table assembly sub-goals? — Align legs → Screw bolts → Tighten.
Table 5: Coverage of the nine primary embodied reasoning dimensions across mainstream benchmarks and our proposed dataset.
Primary dimensions COSMOS ERQA Phyx Where2Place VSI-Bench BLINK RoboSpatial EmbSpatial
Physical & Causal Reasoning
Perception & Object Grounding
Quantitative & Numerical Reasoning
Spatial & Geometric Reasoning
Temporal & Sequential Reasoning
Affordance & Function Reasoning
Multi-Object & Scene Consistency
Scene & Action Understanding
Decision & Task Planning